the sum of a non zero number and its square is equal to three times the number. what is the number?
Answer by checkley77(12844)

x+x^2=3x
x^2+x-3x=0
x^2-2x=0
x(x-2)=0
x-2=0
x=2 ans.
Proof:
2+2^2=3*2
2+4=6
6=6